OpenStudy (anonymous):

y = −x + 2 y = 3x + 1 Which description best describes the solution to the system of equations?

OpenStudy (anonymous):

Do you have answer choices?

OpenStudy (anonymous):

Line y = −x + 2 intersects line y = 3x + 1. Lines y = −x + 2 and y = 3x + 1 intersect the x-axis. Lines y = −x + 2 and y = 3x + 1 intersect the y-axis. Line y = −x + 2 intersects the origin.

OpenStudy (anonymous):

Okay, so what are two ways to solve a system of equations?